Use Cauchy’s Integral formulas to evaluate the following integral along the


indicated closed contours.



z


2+3z+2i


z


2+3z−4


dz;



C



(a) |z| = 2 (b) |z + 5| =


3


2

1
Expert's answer
2022-01-03T17:17:20-0500

Cauchy’s integral formula:

"f(z_0)=\frac{1}{2\pi i}\int_C \frac{f(z)}{z-z_0}dz"

 for any z0 inside C


we have:

"\int_C \frac{z^2+3z+2i}{z^2+3z-4}dz"


"z^2+3z-4=0"

"z=\frac{-3\pm \sqrt{9+16}}{2}"

so,

"z_0=-4,1"

"\frac{z^2+3z+2i}{z^2+3z-4}=\frac{z^2+3z+2i}{(z+4)(z-1)}"

a)

for  |z| = 2:

"z_0=1" is inside C

then:

"f(z)=\frac{z^2+3z+2i}{z+4}"


"\int_C \frac{z^2+3z+2i}{z^2+3z-4}dz=2\pi if(1)=2\pi i(4+2i)/5=\pi(-4 +8i)/5"


b)

for  |z + 5| = 3/2:

"z_0=-4" is inside C

then:

"f(z)=\frac{z^2+3z+2i}{z-1}"


"\int_C \frac{z^2+3z+2i}{z^2+3z-4}dz=2\pi if(-4)=-2\pi i(4+2i)/5=\pi(4 -8i)/5"